File not found error^{∗}

Frank Mittelbach

August 19, 1996

1 Introduction

When L^{A}TEX 2ε is unable to find a file it will ask for an alternative file name. However, sometimes the problem is only noticed by TEX, and in that case TEX insists on getting a valid file name; any other attempt to leave this error loop will fail.^{1} Many users try to respond in the same way as to normal error messages, e.g. by typing ⟨return⟩, or s or x, but TEX will interpret this as a file name and will ask again.

To provide a graceful exit out of this loop, we define a number of files which emulate the normal behavior of TEX in the error loop as far as possible.

After installing these files the user can respond with h, s, e, x, and on some systems also with ⟨return⟩ to TEX’s missing file name question.

∗This file has version v1.0d last revised 1994/12/01

1On some systems, TEX accepts a special character denoting the end of file to return from this loop, e.g. Control-D on UNIX or Control-Z on DOS.